OpenStudy (anonymous):

An apple falls out of a tree from a height of 2.3 m. What is the impact speed of the apple? thanks c:

OpenStudy (abhisar):

Hey @jojolovesbb !

OpenStudy (anonymous):

hello! @Abhisar

OpenStudy (abhisar):

You can use the equation \(v^2=u^2+2\times a\times s\) , Where, v= impact speed, u= initial speed (0) a = acceleration due to gravity (9.8) s = Height of the object (2.3 m in this case.)

OpenStudy (anonymous):

so do i just plug everything together? or what should i do? im not very good in physics unfortunately

OpenStudy (abhisar):

Yes, plug in everything and find the value of V

OpenStudy (abhisar):

Can you do it?

OpenStudy (anonymous):

no i dont think so :( @Abhisar

OpenStudy (abhisar):

Set up the equation like this, \(\sf V^2 = 0 + 2\times 9.8 \times 2.3 \)

OpenStudy (anonymous):

Need help Abster? c;'

OpenStudy (anonymous):

Or you got this one?

OpenStudy (abhisar):

Sure cap, you can take over if you want (-:

OpenStudy (anonymous):

I just pmed her to get back in here

OpenStudy (anonymous):

there she is c;'

OpenStudy (anonymous):

hi c:

OpenStudy (anonymous):

como estas linda? ^.^ te voy ayudar ok...

OpenStudy (anonymous):

okay muchas gracias!

OpenStudy (anonymous):

@Abhisar she doesn't understand cause she's spanish

OpenStudy (anonymous):

nono, i understand english pretty well too lol

OpenStudy (anonymous):

me on the other hand ~cough~ im a Romeo in the Hispanic word ~whips bangs back~

OpenStudy (anonymous):

world* ;-; stupid typos

OpenStudy (anonymous):

haha well hello there c:

OpenStudy (anonymous):

@jojolovesbb do you know the mass of the apple yet?

OpenStudy (anonymous):

no, it makes the equation sort of incomplete

OpenStudy (anonymous):

@Abhisar any ideas?

OpenStudy (anonymous):

i just typed the question as it was shown, i tried to search for any extras and i couldnt find any unfortunately so i guess im gunna have to solve it like this

OpenStudy (anonymous):

@jojolovesbb im trying to think i haven't giving up yet

OpenStudy (anonymous):

okay :o but yeah its missing some information, who the heck came up with this stuff??

OpenStudy (anonymous):

^ no one knows

OpenStudy (anonymous):

yeah :/ oh well lol

OpenStudy (anonymous):

They could of at least said the time of flight

OpenStudy (anonymous):

Im guessing the apple weighs 1 pound

OpenStudy (anonymous):

WAIT @jojolovesbb I got the answer ^_^

OpenStudy (anonymous):

what is it? c:

OpenStudy (anonymous):

http://prntscr.com/7frwg5

OpenStudy (anonymous):

Enough info?

OpenStudy (anonymous):

yay thank you so much!

OpenStudy (anonymous):

np ^_^ have a nice day

OpenStudy (anonymous):

lol eres muy linda senoria ^_^

OpenStudy (anonymous):

senorita*

OpenStudy (anonymous):

aww gracias, usted tambien! :3

OpenStudy (anonymous):

gracias ^.^

OpenStudy (anonymous):

oye mandame un mensaje cuando puedes mi amor c;'